# A Methodology and System For Big-Thick Data Collection

### v2

@@ Abstract
Pervasive sensors have become an essential component in research for gathering real-world data. Nevertheless, current studies on data collecting predominantly concentrate on the objective perception of reality, frequently disregarding the incorporation of subjective human contributions. We introduce an approach and system tailored for gathering big-thick data, a fusion of extensive sensor data (big data) with qualitative human feedback (thick data). Big-thick data enables effective two-way collaboration between humans and machines by supplying data that enables machine learning from human behaviour and activities, as well as data that captures the human interpretation of their actions. The system comprises of three key components: a) a tool for collecting sensor data and user feedback, b) planning and monitoring components for experiment planning and execution monitoring respectively, and c) a machine-learning component that utilizes the data to enhance human-machine interaction
@@ 1 Introduction
Current research on context data collection primarily focuses on context interruptions [H-2017-Mishra], capturing of user attention [mehrotra2016my], and enhancement of question response rates [H-2021-Sun]. However, the current quality of data does not scale to the requirements of big-thick data. A major challenge arises from the fact that the machine necessitates posing a high number of questions to humans, which usually results in low-quality answers [H-2015-Mehrotra.a]. Some approaches depend on fixed or random schedules, which may not align well with participants’ availability or inclination to respond [KD-2021-Zhang-putting], thereby reducing the effectiveness of human-AI collaboration.
In light of the above, our research objective in this paper is to develop a system that enables collection of high quality Big-Thick data, while ensuring minimal disturbance from the AI system. Our aim is to collect Big-Thick data from both an objective and a subjective perspective, and to monitor this process. For data collection, we strive to offer participants flexibility. We integrate a Machine Learning component into the system, which allows the AI to learn about each participant and re- schedule to send context questions at the time when they are willing / available to answer, thereby bolstering the collaboration between humans and AI. For monitoring, our system enables participants to visualize their collected data via a dashboard, thereby providing feedback on their lifestyle and behavior.
The system developed according to the methodology described in this paper, called iLog [KD-2014-PERCOM], has been used extensively in various data collection experiments and studies on human behavior. The first important data collection experiment, called SU2 (Smart University two), collected data about students at the University of Trento, Italy, which data was used in [KD-2021-Zhang-putting] to predict individuals’ behaviors, and in [giunchiglia2018mobile] to show how social media usage can impact students academic performance. A second major data collection was organized in two experiments i.e. DIV1 (Diversity one) and DIV2 (Diversity two) [giunchiglia2021worldwide], to collect data about student behavior, mood, and food habits in 8 different countries [assi2023complex, nanchen2023keep, schelenz2024transparency].
@@ 2 Related Work
To better understand the collected data, visualization can be employed. It effectively supports data exploration, insight communication, and data model improvement / understanding, thereby enabling human-AI collaboration. For example in [samek2017explainable], predictions of a deep learning model are explained through visualization We leverage visualization not only to gain insights into the collected Big-Thick data but also to monitor the progress of the experiment and the overall health of the system, thereby enhancing human-AI collaboration in real-time.
@@ 3 Logical Architecture
Figure 1 depicts the overall system architecture; The light blue rectangles denote the primary components, the calendar and dashboard, with which individuals interact. Dark blue cylinders represent two databases, the first, termed the short term memory (STM) database, stores the plans from both researchers and participants, while the second, named Long Term Memory, the plan execution outcomes from participants. The orange rectangle, termed the scheduler, is the machine learning module that modifies the plan based on participants’ data. The humanoid figures symbolize two roles, i.e., researchers and participants. A researcher is one seeking to gather context information from participants, while a participant is one that imparts their knowledge through the data collection app. The arrow lines illustrate the interactions between the different components, and the mobile phone serves as the tool for collecting data and enables the participant to receive context questions.
The Big-Thick data collection process starts with the researcher’s planning process. He uses the Planner component, in the Calendar module, to plan context questions and schedule sensor data collection. This plan is subsequently stored in the STM database as the scheduled actions which are sent to the data collection application at the set time, where participants can provide answers. Participants can modify the plan through the re-planner component in the Calendar module. Data collected during the execution of the plan is archived in the Long Term Memory database. Plan execution results are sent to the STM database. Furthermore, the scheduler component learns and determines the right time to ask context questions based on the personal context and re-plan modification received from a specific participant. It then sends the predicted time to the STM database to inform the Planning process, ensuring that the system can adapt and respond to changes effectively.
Subsequently, the dashboard provides a means through which both the researcher and participant can view the collected data. They can navigate through, and filter the data thus gaining insights from the collected data. This not only helps the participant become aware of their lifestyle leading to positive improvement (since his behaviour is shown in the data) but also enables him understand how he is fairing in terms of data collected as compared to other participants or the limits set by the researcher. On the other hand, the researcher will gain an overall understanding of the experiment data, including issues which may have occurred and thus mitigate them for the smooth running of the experiment.
@@ 4 Representing Context
The goal of asking context questions or collecting sensor data is to understand the personal context annotation from a participant. The annotated context used in this study is defined in [KD-2017-PERCOM, KD-1993-giunchiglia].
Let’s consider an afternoon at a student’s place where he is having a discussion with his friend. Figure 2 shows this scenario as a knowledge graph, representing the personal context of the student. Each node represents an entity, e.g., person and room, with their respective attributes with values; for instance, attributes of ME (whose context we are describing) are $Class$, $Name$, $Mood$, Notification time and Answer time with the corresponding values as shown in Figure 2. Edges represent relations between entities, e.g., Sitting room is $PartOf$ $Home$, whereas; Peter (person) and the dining table (Table) which are both $in$ the sitting room which $HasActivity$ of a $discussion$ taking place.
@@
  • WI is the internal context generated from the question "What mood are you In??". As shown in Figure 2, it is the emotional state (mood) of me.
@@ 5 Monitoring the Plan Execution
The dashboard, as a visualization component provides flexibility to both the researcher and the participants in monitoring the execution of the experiment plan. It allows them to view and explore the collected big-thick data at any given time. The dashboard helps them gain insights from the data as explained below.
Researcher: The researcher designs an experiment to determine which data to collect and when and how to collect the data. Through the dashboard, he is able to view each participant’s data including the number of questions answered and sensor data collected per participant. He can also compare the data collected between different participants and even filter it using options to get those with the least or most contributions. Not only does it facilitate him to gain insights in the collected data, but it also helps in the monitoring of the experiment progress to ensure that everything goes as planned. Details of the plan execution process including questions sent and data collected (answers and sensors) are displayed on graphs which in turn help in understanding if the plan is progressing as expected or if there is need for mitigation in case of a problem thus ensuring that quality data is collected.
Participant: The participant registers to take part in an experiment and installs the data collection tool on a smart device for the data to be collected. Through the dashboard, participants can navigate through and explore their data to discover insights about themselves, for example, they can learn about their lifestyle depending on the nature and questions of the experiment, e.g., if they eat a lot of snacks, or spend a lot of time in a given place, all of which could point out a negative lifestyle thus prompting them to change to better their life. They are also able to compare how they are faring against other participants in terms of data collected in the experiment. Figure 3 shows such an example, where one participant (green line) is compared to two others (orange and blue lines). The y-axis shows the number of questions answered whereas the x-axis shows the different experiment days. Such comparison prompts the participants improve on their contribution in the experiment.
Figure 3: Comparison of a participant’s answers (green line) with two other participants (blue and orange lines)
@@ 6 Scheduling
The scheduler is a machine learning component which learns and modifies the experiment plan by using a combination of the answers given by a participant and the modifications they make in the re-planner. Let’s explain this with an example;
Given that our participants are students, we try to refrain from sending questions while they are studying. To address this, we use the scheduler component to predict the periods when the participants are likely to be in class, thereby avoiding any disruption to their academic activities. We chose to binary encode study activities. Specifically, studying alone or with others and attending a classroom lecture are encoded as 1, and all other activities as 0. This information was collected from time diaries questionnaire with the question: “What are you doing?”. Other factors can be used to train the machine learning classifiers; including social demographics (i.e., gender, degree, and department), time (i.e., hour and weekday), situational context (i.e., location and interacting individuals), personality traits [donnellan2006mini], human values [schwartz2012overview], and mood.
@@ 7 Conclusion
In this paper, we have presented a novel system for Big-Thick data collection, designed to enhance user experience and improve the quality and quantity of collected data. We have fostered a meaningful collaboration between humans and AI by granting participants the flexibility to choose their response times for context questions and sensor collections. Our system incorporates a visualization component, allowing both researchers and participants to monitor the experiment’s progress and gain insights from the data. Future work will focus on evaluating the system in real-world scenarios, comparing it with existing methods, and addressing potential ethical and privacy implications.

### v3

@@ Abstract
Pervasive sensors have become essential in research for gathering real-world data. However, current studies often focus solely on objective data, neglecting subjective human contributions. We introduce an approach and system for collecting big-thick data, combining extensive sensor data (big data) with qualitative human feedback (thick data). This fusion enables effective collaboration between humans and machines, allowing machine learning to benefit from human behavior and interpretations. Emphasizing data quality, our system incorporates continuous monitoring and adaptive learning mechanisms to optimize data collection timing and context, ensuring relevance, accuracy, and reliability. The system comprises three key components: a) a tool for collecting sensor data and user feedback, b) components for experiment planning and execution monitoring, and c) a machine-learning component that enhances human-machine interaction.
@@ 1 Introduction
Current research on context data collection primarily focuses on context interruptions [H-2017-Mishra], capturing of user attention [mehrotra2016my], and enhancement of question response rates [H-2021-Sun]. However, the current data quality does not meet the demands of big-thick data. A significant challenge arises from machines needing to pose a high volume of questions to humans, often resulting in low-quality responses [H-2015-Mehrotra.a]. Certain methods also rely on fixed or random schedules that may not align with participants’ availability or willingness to respond [KD-2021-Zhang-putting], thereby affecting the quality of the data.
In light of the above, our research objective in this paper is to develop a system that enables collection of high quality Big-Thick data, while ensuring minimal disturbance from the AI system. Our aim is to collect quality Big-Thick data from both an objective and a subjective perspective, while monitoring this process. We offer participants flexibility by incorporating a Machine Learning component which adaptively schedules context questions based on participants’ availability and willingness to respond as a means to improve data quality. For monitoring, our system enables visualizing the collected data via a dashboard, thereby providing real-time feedback on the quality for example, missing sensor and question data.
The system developed according to the methodology in this paper is an improvement and extension of iLog [KD-2014-PERCOM], which has been widely used in various data collection experiments and studies on human behavior. The first experiment, SU2 (Smart University Two), gathered data on students at the University of Trento, Italy. This data was used in [KD-2021-Zhang-putting] to predict individuals’ behaviors, and in [giunchiglia2018mobile] to examine the impact of social media usage on students’ academic performance. Another major data collection involved two experiments, DIV1 (Diversity One) and DIV2 (Diversity Two) [giunchiglia2021worldwide], which collected data on student behavior, mood, and food habits across eight countries. The extensions presented in this paper are motivated by the goal of having a fully integrated human-in-the-loop human machine interaction [giunchiglia2022contextmodelpersonaldata, bontempelli2022lifelongpersonalcontextrecognition].
@@ 2 Related Work
To better understand the collected data, visualization can be employed. It effectively supports data exploration, insight communication, and data model improvement / understanding, for quality data. For example in [samek2017explainable], predictions of a deep learning model are explained through visualization We leverage visualization not only to gain insights into the collected Big-Thick data but also to monitor the progress of the experiment for quality data collection.
@@ 3 Logical Architecture
Figure 1 depicts the overall system architecture; The light blue rectangles denote the primary components, the calendar and dashboard, with which individuals interact. Dark blue cylinders represent two databases, the first, termed the short term memory (STM) database, stores the plans from both researchers and participants, while the second, named Long Term Memory, the plan execution outcomes from participants. The orange rectangle represents the scheduler, a machine learning module that modifies plans based on participants’ data. The humanoid figures symbolize two roles, i.e., researchers and participants. A researcher is one seeking to gather context information from participants, while a participant is one that imparts their knowledge through the data collection app. The arrow lines illustrate the interactions between the different components, and the mobile phone serves as the tool for collecting data and enables the participant to receive context questions.
The Big-Thick data collection process begins with the researcher’s planning. Using the Planner component within the Calendar module, the researcher schedules context questions and sensor data collection. This plan is stored in the STM database as scheduled actions, which are then sent to the data collection application at the designated times for participants to provide responses. Participants can modify the plan through the re-planner component in the Calendar module. Data collected during the execution of the plan is archived in the Long Term Memory database, while plan execution results are sent to the STM database. Furthermore, the scheduler component learns the optimal times to ask context questions based on the personal context and re-plan modifications from participants. It then sends these predicted times to the STM database to inform the planning process, ensuring the system adapts and responds effectively to changes.
Subsequently, the dashboard provides a means for both the researcher and participant to view and interact with the collected data. They can navigate through and filter the data, gaining valuable insights. For participants, this awareness of their lifestyle, as reflected in the data, can lead to positive behavioral improvements. Participants can also see how their data compares to that of other participants or to the limits set by the researcher.The dashboard provides researchers with a comprehensive overview of the experimental data, including any encountered problems. This enables them to promptly address these difficulties to ensure the seamless execution of the experiment for accurate data collection..
@@ 4 Representing Context
The goal of asking context questions or collecting sensor data is to understand the personal context annotation from a participant. The annotated context used in this study is defined in [KD-2017-PERCOM, KD-1993-giunchiglia]. As an example, let’s consider an afternoon at a student’s place where he is in a discussion with his friend. Figure 2 shows this scenario as a knowledge graph, representing the personal context of the student. Each node represents an entity, e.g., person and room, with their respective attributes with values; for instance, attributes of ME (whose context we are describing) are $Class$, $Name$, $Mood$, Notification time and Answer time with the corresponding values as shown in Figure 2. Edges represent relations between entities, e.g., Sitting room is $PartOf$ $Home$, whereas; Peter (person) and the dining table (Table) which are both $in$ the sitting room which $HasActivity$ of a $discussion$ taking place.
@@
  • WI is the internal context generated from the question "What mood are you In?". As shown in Figure 2, it is the emotional state (mood) of me.
@@ 5 Monitoring the Plan Execution
The dashboard, as a visualization component offers flexibility to both the researcher and the participants in monitoring the execution of the experiment plan. It allows them to view and explore the collected big-thick data at any given time. The dashboard helps them gain insights from the data as explained below.
Researcher: The researcher designs an experiment to determine which data to collect and when and how to collect the data. Through the dashboard, he is able to view each participant’s data including the number of questions answered and sensor data collected per participant. He can also compare the data collected between different participants and even filter it using options to get those with the least or most contributions. Not only does it facilitate him to gain insights in the collected data, but it also helps in the monitoring of the experiment progress to ensure that everything goes as planned. Details of the plan execution process including questions sent and data collected (answers and sensors) are displayed on graphs which in turn help in understanding if the plan is progressing as expected or if mitigation is needed, ensuring the collection of high-quality data.
Participant: The participant registers to take part in an experiment and installs the data collection tool on a smart device. Through the dashboard, participants can navigate and explore their data to gain insights about themselves. For example, depending on the experiment’s nature and questions, they can learn about their lifestyle, such as if they eat a lot of snacks or spend excessive time in a particular location. Such insights can highlight negative behaviors and prompt positive lifestyle changes. Participants can also compare their data against other participants to see how they fare in terms of data collected in the experiment. Figure 3 shows such an example, where one participant (green line) is compared to three others (orange, red and blue lines). The y-axis shows the number of questions answered whereas the x-axis shows the different experiment days. Such comparison prompts the participants improve on their contribution in the experiment.
Figure 3: Comparison of a participant’s answers (green line) with three other participants (blue, red and orange lines)
@@ 6 Scheduling
The scheduler is a machine learning component which learns and modifies the experiment plan by using a combination of the answers given by a participant and the modifications they make in the re-planner. Let’s explain this with an example: Given that our participants are students, we try to refrain from sending questions while they are studying. To address this, we use the scheduler component to predict the periods when the participants are likely to be in class, thereby avoiding any disruption to their academic activities.
Using data from the WeNet experiment described in [giunchiglia2021worldwide], we trained Random Forests, Decision Trees, Artificial Neural Networks, Logistic Regression, and Gaussian Naive Bayes on comprehensive training and testing sets encompassing all participants (170 students). We chose to binary encode study activities. Specifically, studying alone or with others and attending a classroom lecture are encoded as 1, and all other activities as 0. This information was collected from time diaries questionnaire with the question: “What are you doing?”. The results are shown in Tab 1 with the Random Forest classifier showing the highest prediction accuracy of 75.1%. Other features that can be used for training include; social demographics (e.g gender, department), time (e.g hour, weekday), situational context (e.g location, interacting individuals), personality traits [donnellan2006mini], and mood.
Table 1: Prediction results of different machine learning classifiers on Wenet Data.
Classifier Accuracy Kappa Precision Recall F1 score AUC
Random Forest 0.7510 0.3832 0.6903 0.4491 0.5442 0.8146
Decision Tree 0.7234 0.3338 0.6091 0.4587 0.5233 0.6821
Artificial Neural Networks 0.7209 0.3702 0.5781 0.5799 0.5790 0.7762
Logistic Regression 0.4557 0.3536 0.4892 0.0740 0.1286 0.6705
Gaussian Naive Bayes 0.6295 0.2388 0.4571 0.6366 0.5321 0.6676
@@ 7 Conclusion
In this paper, we have presented a novel system for quality Big-Thick data collection, designed to enhance user experience and improve the quality and quantity of collected data. We have fostered a meaningful collaboration between humans and AI by granting participants the flexibility to choose their response times for context questions and sensor collections. Our system incorporates a visualization component, allowing both researchers and participants to monitor the experiment’s progress and gain insights from the data. Future work will focus on evaluating the system in real-world scenarios, comparing it with existing methods, and addressing potential ethical and privacy implications.
Acknowledgment
The research by Fausto, Ivan, and Leonardo were funded by the European Union’s Horizon 2020 FET Proactive project “WeNet – The Internet of us”, grant agreement No 823783. The work by Haonan received funding from the China Scholarships Council (No.202107820038).